OpenStudy (anonymous):

find the slope and y-intercept of the line: g(x)= -3x-17.5

hartnn (hartnn):

y = g(x) = mx+c has slope of m and y-intercept of c so what about g(x)= -3x-17.5 ?

OpenStudy (anonymous):

Remember y=mx+b Everything falls in place to find the slope & y-intercept there. Y intercept = b and slope=m

OpenStudy (anonymous):

oh so -3 is the slope and the y-intercept is -17.5

hartnn (hartnn):

yup

OpenStudy (anonymous):

yes, correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Correct. You answered your own question (:

OpenStudy (anonymous):

who here is good at radical equations? http://openstudy.com/study#/updates/5060e298e4b02e13941115b7